OpenStudy (handsomechickenlover):

Simlify -4.5[2x-(0.5-1.4)] a. -9x+0.9 b.-9x-0.9 c.-9x-4.05 d. 9x-4.05

OpenStudy (handsomechickenlover):

I need some help with this. Will medal!

OpenStudy (handsomechickenlover):

I got -4.5[2x-(-9)] but need help simplifying it better.

OpenStudy (handsomechickenlover):

SOSOSOSOSOSOSOSOS

OpenStudy (handsomechickenlover):

HEEELLLLPPPPPPPP

OpenStudy (handsomechickenlover):

WILL Medal and fan!!!!!!!!!!!!!!!!!

OpenStudy (khalid3166):

C

OpenStudy (handsomechickenlover):

Thanks

OpenStudy (khalid3166):

np

OpenStudy (handsomechickenlover):

Could someone show me ow to do it now?

OpenStudy (danjs):

if you see parenthesis or brackets, start there with the inside most thing

OpenStudy (danjs):

-4.5*[ 2x- (-0.9) ] parenthesis or brackets are the most priorty

OpenStudy (danjs):

minus a negative is the same as addition a-(-b) = a+b -4.5*[2x + 0.9]

OpenStudy (danjs):

if you have a thing multiplied by a quantity like that, you can 'distribute' the outside term into all the ( ) terms a*(b+c) = a*b + a*c -4.5*[2x + 0.9] = (-4.5)*2x + (-4.5)*0.9 then you left with addition and multiplication muiltiply before adding,

OpenStudy (handsomechickenlover):

Thanks